The total number of students in Year 8 at a high school is 180. There are 24 more girls than boys. How many boys are there??
Answer by Maths68(1474)


Let
Boys=b
Girls=g

Total Number of Students = 180
b+g=180..................(1)
There are 24 more girls than boys
g=24+b....................(2)
Substitute the value of g from (1) to (2)
b+24+b=180
b+24+b=180
2b=180-24
2b=156
2b/2=156/2
b=78
Put the value of b in (2)
g=24+b
g=24+78
g=102
Boys=b=78
Girls=g=102